OpenStudy (anonymous):

Find the midpoint of the segment with the given endpoints. D(1, 2) and E(–3, 6)

OpenStudy (anonymous):

\[\huge Midpoints = (\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2})\]

OpenStudy (anonymous):

\[(x_1, y_1) = (1, 2)\] \[(x_2, y_2) = (-3. 6)\]

OpenStudy (anonymous):

Use the midpoint formula: Add the x coordinates, and the y coordinates and divide by 2. 1 + -3 = -2 Divide by 2: -1 2 + 6 = 8 Divide by 2: 4 so the midpoint is: (-1,4)

OpenStudy (anonymous):

No, its (-1, 4)..